OpenStudy (anonymous):

use the elimination method to solve the system of equations. 2x+5y=-3 x+5/4y=-1/2

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

so look at all the constants in front of x and y, there is one that has 1 as a constant in the second equation, we can use it to get rid of 2x in the 1st equation. But what do we need to multiply x with (in the second equation) then add to the 1st to get rid of that 2x?

OpenStudy (anonymous):

im confused with this 1 @ybarrap

OpenStudy (ybarrap):

what happens if we multiply the 2nd row by -2, what do we got then?

OpenStudy (ybarrap):

just multiply -2 times row 2 and write the results, I promise that this will be easy

OpenStudy (ybarrap):

-2(row2) = -2 * (x+5/4y )= -2*( -1/2) What doe this look like after multiplying out?

OpenStudy (anonymous):

idk im lost @ybarrap

OpenStudy (ybarrap):

-2 * (x+5/4y )= -2*( -1/2) means -2x - (5/2) y = 1 Do you see how this was done?

OpenStudy (anonymous):

i think so u took 2 from the 5/4 and -1/2 made 1 @ybarrap

OpenStudy (ybarrap):

no, -2 * (5/4) = -5/2 and -2 * (-1/2) = 1 That's how I got each of these values

OpenStudy (ybarrap):

Any problems with this part? If you do, it'll be hard to go on

OpenStudy (anonymous):

continue plz @ybarrap

OpenStudy (ybarrap):

You'll need to add this to row 1, replacing row 1 with the result. This will need to be my final advice for you.